### v1

# Spatially covariant gravity with two degrees of freedom in the presence of an auxiliary scalar field: Hamiltonian analysis

Abstract
A class of gravity theories respecting spatial covariance and in the presence of non-dynamical auxiliary scalar fields with only spatial derivatives is investigated. Generally, without higher temporal derivatives in the metric sector, there are 3 degrees of freedom (DOFs) propagating due to the breaking of general covariance. Through a Hamiltonian constraint analysis, we examine the conditions to eliminate the scalar DOF such that only 2 DOFs, which correspond the tensorial gravitational waves in a homogeneous and isotropic background, are propagating. We find that two conditions are needed, each of which can eliminate half degree of freedom. The second condition can be further classified into two cases according to its effect on the Dirac matrix. We also apply the formal conditions to a polynomial-type Lagrangian as a concrete example, in which all the monomials are spatially covariant scalars containing two derivatives. Our results are consistent with the previous analysis based on the perturbative method.
I Introduction
With the detection of gravitational waves, a novel tool has emerged for testing theories of gravity. A fundamental question arises: Is general relativity the only theory of gravity that propagates gravitational waves with precisely two degrees of freedom (2DOFs)? This question is partially addressed by Lovelock’s theorem [1, 2], which states that in four dimensions, if one requires spacetime covariance and second-order equations of motion, general relativity is indeed the unique theory that propagates gravitational waves with two degrees of freedom. This also implies that, by violating the assumptions of Lovelock’s theorem, other gravitational theories propagating two degrees of freedom can, in principle, exist.
Gravitational theories with 2DOFs can be traced back to the Cuscuton theory [3], which has been extensively studied [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], and extended in the framework of scalar-tensor theories with higher-order derivatives [17, 18]. Such theories have also been discussed as special cases of Hořava gravity [19, 20, 21, 22]. The connection between 2DOFs and spacetime symmetries has been investigated in [23, 24, 25, 26].
Another class of theories propagating 2DOFs was proposed in [27], known as minimally modified gravity (MMG) [^1] See also the minimal theory of massive gravity studied in [28, 29, 30, 31, 32, 33].. The so-called type-II MMG [34, 35] includes the original Cuscuton theory as a special case. MMG has been extensively studied [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], and has also been generalized by introducing auxiliary constraints in phase space [52, 53].
A general framework for spatially covariant gravity (SCG), which respects only spatial covariance, was proposed in [54, 55]. Due to the breaking of full spacetime covariance, SCG generally propagates a scalar degree of freedom in addition to the 2DOFs corresponding to gravitational waves. Thus, it provides an alternative approach to constructing scalar-tensor theories [56, 57, 58, 59]. SCG includes Hořava gravity [60, 61, 62] and the effective field theory of inflation/dark energy [63, 64, 65, 66, 67] as special cases. It has been further generalized to include a dynamical lapse function [68, 69], nonmetricity [70], parity violation [71], and multiple scalar modes [72]. SCG has also been applied in cosmology and gravitational wave studies [73, 74, 75, 76, 77]. Subclasses of SCG that propagate only 2DOFs were explored in [78, 79], and in the presence of a dynamical lapse function in [80]. The explicit Lagrangian found in [78] has been employed in cosmological studies [81, 82, 83, 10, 84].
In this work, we generalize the SCG framework to include an auxiliary scalar field and determine the conditions under which the theory propagates only 2DOFs. The idea of spatially covariant gravity with an auxiliary scalar field was first introduced in [85], originally motivated by generally covariant scalar-tensor theories in which the scalar field has a spacelike gradient [^2] However, we emphasize that these are entirely different theories. See Appendix A for details.. The presence of an auxiliary scalar field introduces new and novel possibilities for constructing 2DOF theories.
In [86], a perturbative analysis similar to [17, 69, 79] was performed to derive necessary conditions for the absence of scalar modes at linear order in perturbations around a cosmological background. Nevertheless, a Hamiltonian constraint analysis is still required to determine the necessary and sufficient conditions for the theory to propagate only 2DOFs at the nonlinear level. This paper is devoted to this task.
The paper is organized as follows. In Sec. II, we describe our model of spatially covariant gravity with an auxiliary scalar field. We find that new terms in the Lagrangian are allowed due to the inclusion of the auxiliary scalar. In Sec. III, we derive the Hamiltonian formalism and show that, in general, a scalar mode propagates unless additional constraints are imposed. In Sec. IV, we perform a degeneracy analysis and derive the conditions under which the theory propagates only 2DOFs. In Sec. V, we apply our formalism to a concrete model, where the Lagrangian is built from SCG monomials with $d=2$. Finally, in Sec. VI, we summarize our results.
II Spatially covariant gravity with an auxiliary scalar field
Our starting point is the action
$S=\int\mathrm{d}t\mathrm{d}^{3}xN\sqrt{h}\mathcal{L}\left(\phi,N,h_{ij},{}^{3}\!R_{ij};\mathrm{D}_{i},\pounds_{\bm{n}}\right),$ (1)
where $N$ is the usual lapse function, $h_{ij}$ is the spatial metric on the spacelike hypersurfaces, ${}^{3}\!R_{ij}$ is the spatial Ricci tensor, $\mathrm{D}_{i}$ is the spatially covariant derivative compatible with $h_{ij}$, and £$\pounds_{\bm{n}}$ is the Lie derivative with respect to the normal vector $n^{a}$ of the hypersurfaces. By definition, the action (1) respects only the three-dimensional spatial invariance. We require that the Lie derivative £$\pounds_{\bm{n}}$ acts only on $h_{ij}$, which yields the extrinsic curvature $K_{ij}$ defined by
£$\pounds_{\bm{n}}h_{ij}=2K_{ij}=\frac{1}{N}\left(\dot{h}_{ij}-\pounds_{\vec{N}}h_{ij}\right).$ (2)
As a result, both the lapse function $N$ and the scalar field $\phi$ play the role of auxiliary variables. Note that the shift vector $N^{i}$ should not appear explicitly in the Lagrangian, as it is actually the gauge field of spatial covariance.
Since the Lie derivative enters the Lagrangian only through the extrinsic curvature, we simply introduce $K_{ij}$ to replace the Lie derivative £$\pounds_{\bm{n}}$ in (1), and consider the following action:
$\tilde{S}=\int\mathrm{d}t\mathrm{d}^{3}x\,N\sqrt{h}\mathcal{L}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},K_{ij};\mathrm{D}_{i}\right).$ (3)
According to the Hamiltonian analysis in the next section, the action (3) generally has 3 DOFs. This is simply because $\phi$ has no dynamics; the number of DOFs in (3) is the same as in spatially covariant gravity without the scalar field [54, 55].
In principle, Lie derivatives of the lapse function £$\pounds_{\bm{n}}N$ and of the extrinsic curvature £$\pounds_{\bm{n}}K_{ij}$ can be considered, which would introduce additional degrees of freedom. It has been shown that extra conditions must be imposed in order to eliminate these unwanted degrees of freedom. For example, the theory will propagate an extra DOF (i.e., a total of 4 DOFs) in the case of a dynamical lapse function. Two conditions are required to eliminate such an extra DOF [68]. A similar analysis was carried out to fully eliminate the scalar DOFs [80].
Generally, the Lagrangian in (3) may be nonlinear in $K_{ij}$, which makes the explicit inversion of velocity $\dot{h}_{ij}$ in terms of the momentum $\pi^{ij}$ impossible. This problem can be solved by rewriting (3) in an equivalent form:
$\tilde{S}$ $=\int\mathrm{d}t\mathrm{d}^{3}xN\sqrt{h}\tilde{\mathcal{L}}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},K_{ij};\mathrm{D}_{i}\right)$
$=S\left(\phi,N,h_{ij},{}^{3}\!R_{ij},B_{ij};\mathrm{D}_{i}\right)+\int\mathrm{d}t\mathrm{d}^{3}xN\sqrt{h}\varLambda^{ij}\left(K_{ij}-B_{ij}\right),$ (4)
where the action $S\left(\phi,N,h_{ij},{}^{3}R_{ij},B_{ij};\mathrm{D}_{i}\right)$ is the same as (3) with $K_{ij}$ replaced by $B_{ij}$. This strategy of using a Lagrange multiplier to linearize the velocities was employed in [68, 78]. The equations of motion for $\varLambda^{ij}$ ensure that the auxiliary field $B_{ij}$ is exactly equal to $K_{ij}$. Thus, the two actions are equivalent, at least at the classical level. By making use of the equation of motion for the auxiliary field $B_{ij}$, we can determine the Lagrange multiplier $\varLambda^{ij}$ as
$\varLambda^{ij}=\frac{1}{N\sqrt{h}}\frac{\delta S}{\delta B_{ij}}.$ (5)
Finally, by plugging the solution for $\varLambda^{ij}$ into the action (4), we can recast it in the form
$\tilde{S}=S\left(\phi,N,h_{ij},{}^{3}\!R_{ij},B_{ij};\mathrm{D}_{i}\right)+\int\mathrm{d}t\mathrm{d}^{3}x\frac{\delta S}{\delta B_{ij}}\left(K_{ij}-B_{ij}\right).$ (6)
In the following, we will use (6) as our starting point for the Hamiltonian analysis.
III Hamiltonian formalism
III.1 Hamiltonian and primary constraints
In the action (6), the set of variables is
$\varPhi^{I}\coloneqq\left\{N^{i},\phi,B_{ij},N,h_{ij}\right\},$ (7)
and the corresponding set of conjugate momenta is denoted by
$\varPi_{I}\coloneqq\left\{\pi_{i},p,p^{ij},\pi,\pi^{ij}\right\}.$ (8)
The conjugate momenta are defined as
$\varPi_{I}$ $\coloneqq\frac{\delta\tilde{S}}{\delta\dot{\Phi}^{I}}.$ (9)
Since all velocities $\dot{\varPhi}^{I}$ in the Lagrangian (6) appear linearly, the conjugate momenta automatically yield constraints among the phase space variables. Indeed, the conjugate momenta are
$\left\{\pi_{i},p,p^{ij},\pi,\pi^{ij}\right\}$ $=$ $\left\{\frac{\delta\tilde{S}}{\delta\dot{N}^{i}},\frac{\delta\tilde{S}}{\delta\dot{\phi}},\frac{\delta\tilde{S}}{\delta\dot{B}_{ij}},\frac{\delta\tilde{S}}{\delta\dot{N}},\frac{\delta\tilde{S}}{\delta\dot{h}_{ij}}\right\}$ (10)
$=$ $\left\{0^{i},0,0_{ij},0,\frac{1}{2N}\frac{\delta S}{\delta B_{ij}}\right\},$
which lead to the primary constraints
$\pi_{i}\approx 0,\quad p\approx 0,\quad p^{ij}\approx 0,\quad\pi\approx 0,\quad\tilde{\pi}^{ij}\equiv\pi^{ij}-\frac{1}{2N}\frac{\delta S}{\delta B_{ij}}\approx 0.$ (11)
Here and throughout this work, the symbol $\approx$” denotes “weak equality”, i.e., an equality that holds only when the constraints are satisfied. By contrast, “$=$” denotes strong equality, which holds regardless of whether constraints are imposed.
We denote the set of primary constraints as
$\varphi^{I}\coloneqq\left\{\pi_{i},p,p^{ij},\pi,\tilde{\pi}^{ij}\right\},$ (12)
and define the subspace of the phase space where the primary constraints are satisfied as $\varGamma_{\mathrm{P}}$, which can be regarded as a hypersurface in the original phase space $\left\{\varPhi^{I},\varPi_{I}\right\}$.
The canonical Hamiltonian is obtained via the Legendre transformation:
$H_{\mathrm{C}}$ $=\int\mathrm{d}^{3}x\left(\sum_{I}\varPi_{I}\dot{\varPhi}^{I}-N\sqrt{h}\tilde{\mathcal{L}}\right)\approx\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}},$ (13)
where
$\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}}=\int\mathrm{d}^{3}x(NC)+X[\vec{N}],$ (14)
with
$C=2\pi^{ij}B_{ij}-\sqrt{h}\mathcal{L}.$ (15)
For a general “spatial” vector field $\vec{\xi}$, $X[\vec{\xi}]$ is defined as [68]
$X[\vec{\xi}]$$\coloneqq$ $\int\mathrm{d}^{3}x\,\varPi_{I}\pounds_{\vec{\xi}}\varPhi^{I}$ (16)
$=$ $\int\mathrm{d}^{3}x\left(\pi_{i}\pounds_{\vec{\xi}}N^{i}+p\pounds_{\vec{\xi}}\phi+p^{ij}\pounds_{\vec{\xi}}B_{ij}+\pi\pounds_{\vec{\xi}}N+\pi^{ij}\pounds_{\vec{\xi}}h_{ij}\right).$
In the following, we briefly summarize the properties of $X[\vec{\xi}]$ as discussed in [68]. Suppose $Q\left(\vec{x}\right)$ is a general spatial tensor field or density (with spatial indices suppressed for brevity) constructed from the phase space variables $\left\{\varPhi^{I},\varPi_{I}\right\}$, and let $\mathcal{F}$ be a scalar functional defined by
$\mathcal{F}\left[f,\varPhi^{I},\varPi_{I}\right]=\int\mathrm{d}^{3}x\,f(\vec{x})Q(\vec{x}),$ (17)
where $f(\vec{x})$ is a smooth test function. Note that $f(\vec{x})$ may also carry appropriate spatial indices (suppressed here) such that $\mathcal{F}$ is invariant under time-independent spatial diffeomorphisms. The Poisson bracket between $X[\vec{\xi}]$ and $\mathcal{F}$ is given by
$[X[\vec{\xi}],\mathcal{F}]$ $=$ $X\left[[\vec{\xi},\mathcal{F}]\right]+\int\mathrm{d}^{3}x\frac{\delta\mathcal{F}}{\delta f}\pounds_{\vec{\xi}}f$ (18)
$=$ $\int\mathrm{d}^{3}x\,Q(\vec{x})\pounds_{\vec{\xi}}f(\vec{x}).$
In (18), the Poisson bracket $[\mathcal{F},\mathcal{G}]$ is defined as
$[\mathcal{F},\mathcal{G}]\coloneqq\int\mathrm{d}^{3}z\left(\frac{\delta\mathcal{F}}{\delta\varPhi_{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPi^{I}(\vec{z})}-\frac{\delta\mathcal{F}}{\delta\varPi^{I}(\vec{z})}\frac{\delta\mathcal{G}}{\delta\varPhi_{I}(\vec{z})}\right).$ (19)
A key point is that the right-hand side of the second line in (18) vanishes on the constraint surface if $Q(\vec{x})$ is a constraint.
By integrating by parts, $X[\vec{\xi}]$ can be written in a more convenient form:
$X[\vec{\xi}]=\int\mathrm{d}^{3}x\,\xi^{i}\mathcal{C}_{i},$ (20)
where
$\mathcal{C}_{i}$ $=$ $\pi\,\mathrm{D}_{i}N-2\sqrt{h}\mathrm{D}_{j}\left(\frac{1}{\sqrt{h}}\pi_{i}^{j}\right)+p\,\mathrm{D}_{i}\phi+p^{kl}\mathrm{D}_{i}B_{kl}$ (21)
$-2\sqrt{h}\mathrm{D}_{j}\left(\frac{p^{jk}}{\sqrt{h}}B_{ik}\right)+\pi_{j}\mathrm{D}_{i}N^{j}+\sqrt{h}\mathrm{D}_{j}\left(\frac{1}{\sqrt{h}}\pi_{i}N^{j}\right).$
As we shall see later, the constraints $\mathcal{C}_{i}\approx 0$ are secondary constraints associated with the primary constraints $\pi_{i}\approx 0$. Furthermore, the Poisson bracket between $\mathcal{C}_{i}$ and any constraint $Q$ is given by
$\left[X[\vec{\xi}],\mathcal{F}\right]=\int\mathrm{d}^{3}x\mathrm{d}^{3}y\,\xi^{i}(x)f(y)\left[\mathcal{C}_{i}(\vec{x}),Q(\vec{y})\right].$ (22)
It immediately follows from (18) that
$\left[\mathcal{C}_{i}(\vec{x}),Q(\vec{y})\right]\approx 0,\quad\text{for any}\quad Q\approx 0.$ (23)
Due to the presence of primary constraints, the time evolution is governed by the total Hamiltonian, defined as
$H_{\mathrm{T}}\coloneqq H_{\mathrm{C}}+\int\mathrm{d}^{3}y\sum_{I}\lambda_{I}(\vec{y})\varphi^{I}(\vec{y}),$ (24)
where $\lambda_{I}:=\left\{v^{i},v,v_{ij},\lambda,\lambda_{ij}\right\}$ are undetermined Lagrange multipliers. Since the canonical Hamiltonian $H_{\mathrm{C}}$ is only well-defined on the subspace defined by the primary constraints $\varGamma_{\mathrm{P}}$, we may directly use $\left.H_{\mathrm{C}}\right|_{\varGamma_{\mathrm{P}}}$ given in (14) in place of $H_{\mathrm{C}}$ in (13) for subsequent calculations.
III.2 Consistency condition
Constraints must be preserved under time evolution. For the primary constraints, we require that
$\int\mathrm{d}^{3}y\left[\varphi^{I}(\vec{x}),\varphi^{J}(\vec{y})\right]\lambda_{J}(\vec{y})+\left[\varphi^{I}(\vec{x}),H_{\mathrm{C}}\right]\approx 0,$ (25)
which are the so-called consistency conditions for the primary constraints. These conditions may yield further constraints.
Without imposing additional assumptions on the structure of the Lagrangian in (6), the Poisson brackets between the primary constraints have been computed in detail in [85], and are summarized in Appendix B. For later convenience, we rewrite (25) in matrix form:
$\int\mathrm{d}^{3}y\left(\begin{array}[]{ccccc}0&0&0&0&0\\ 0&0&0&0&\left[p(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\\ 0&0&0&0&\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\\ 0&0&0&0&\left[\pi(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\\ 0&\left[\tilde{\pi}^{ij}(\vec{x}),p(\vec{y})\right]&\left[\tilde{\pi}^{ij}(\vec{x}),p^{kl}(\vec{y})\right]&\left[\tilde{\pi}^{ij}(\vec{x}),\pi(\vec{y})\right]&\left[\tilde{\pi}^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\end{array}\right)\left(\begin{array}[]{c}v^{i}\\ v\\ v_{ij}\\ \lambda\\ \lambda_{ij}\end{array}\right)\approx\left(\begin{array}[]{c}-\mathcal{C}_{i}(\vec{x})\\ \frac{\delta S}{\delta\phi(\vec{x})}\\ 0\\ -\mathcal{C}(\vec{x})\\ \left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]\end{array}\right),$ (26)
where $\mathcal{C}_{i}$ is defined in (22). Note that $\mathcal{C}_{i}$ must have vanishing Poisson brackets with all other constraints due to the property (22). In the fourth line of (26), $\mathcal{C}$ is defined by
$\mathcal{C}=2\pi^{ij}(\vec{x})B_{ij}(\vec{x})-\frac{\delta S}{\delta N(\vec{x})},$ (27)
and $\left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]$ in the fifth line is given by
$\left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]=\frac{\delta S}{\delta h_{ij}(\vec{x})}-\frac{1}{N(\vec{x})}\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta h_{kl}(\vec{z})}N(\vec{z})B_{kl}(\vec{z}).$ (28)
According to (129), $\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ in the third line of (26) is proportional to the kinetic matrix
$\frac{\delta^{2}S_{B}}{\delta B_{ij}(\vec{x})\delta B_{kl}(\vec{y})},$ (29)
which we assume to be non-degenerate. This is because $B_{ij}$ is the auxiliary field corresponding to $K_{ij}$. The non-degeneracy of (29) implies that the kinetic term for $\dot{h}_{ij}$ is non-degenerate, which guarantees that the theory includes General Relativity as a limiting case, and in particular, allows the existence of gravitational waves.
As a result, the Lagrange multipliers $\lambda_{ij}$ can be fixed from the third line of (26) as
$\lambda_{ij}=0.$ (30)
Then the second and fourth lines of (26) yield two additional secondary constraints:
$\frac{\delta S}{\delta\phi(\vec{x})}\approx 0,\quad\mathcal{C}(\vec{x})\approx 0.$ (31)
The constraints $\frac{\delta S}{\delta\phi(\vec{x})}$ and $\mathcal{C}(\vec{x})$ do not generate further constraints via consistency conditions, as will become clear in the next section.
Having identified all the primary and secondary constraints, we are now in a position to classify them. According to the terminology of the Dirac-Bergmann algorithm, a constraint that has vanishing Poisson brackets with all other constraints (including itself) is called a first-class constraint; otherwise, it is second-class. The distinction is that each first-class constraint eliminates one degree of freedom, whereas a second-class constraint removes one phase space dimension, i.e., half a degree of freedom.
From the discussion above on $\mathcal{C}_{i}$ and the constraint algebra in Appendix B, the Poisson brackets among all the constraints can be summarized in the so-called Dirac matrix:
$\begin{array}[]{|c|cccccccc|}\hline\cr[\cdot,\cdot]&\pi_{i}(\vec{y})&\mathcal{C}_{i}(\vec{y})&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})&\tilde{\pi}^{kl}(\vec{y})&\mathcal{C}(\vec{y})&\frac{\delta S}{\delta\phi(\vec{y})}\\ \hline\cr\pi_{i}(\vec{x})&0&0&0&0&0&0&0&0\\ \mathcal{C}_{i}(\vec{x})&0&0&0&0&0&0&0&0\\ p(\vec{x})&0&0&0&0&0&X&X&X\\ \pi(\vec{x})&0&0&0&0&0&X&X&X\\ p^{ij}(\vec{x})&0&0&0&0&0&X&X&X\\ \tilde{\pi}^{ij}(\vec{x})&0&0&X&X&X&X&X&X\\ \mathcal{C}(\vec{x})&0&0&X&X&X&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&0&0&X&X&X&X&X&0\\ \hline\cr\end{array},$ (32)
where an “$X$” denotes a generically non-vanishing Poisson bracket. In total, there are 22 constraints, which can be divided into two classes:
$\begin{array}[]{rl}6\text{ first-class: }&\pi_{i},\quad\mathcal{C}_{i},\\ 16\text{ second-class: }&p,\quad p^{ij},\quad\pi,\quad\tilde{\pi}^{ij},\quad\frac{\delta S}{\delta\phi},\quad\mathcal{C}.\end{array}$
The number of degrees of freedom (DOF) is thus
#$\#_{\mathrm{DOF}}$ $=$ $\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (33)
$=$ $\frac{1}{2}(2\times 17-2\times 6-16)$
$=$ 3 3,3 ,
which explicitly shows that a spatially covariant gravity theory where temporal derivatives arise only through the extrinsic curvature has 3 propagating degrees of freedom, regardless of the presence of additional auxiliary field(s).
IV Degeneracy Analysis
To reduce the number of DOFs, additional conditions must be imposed on the Lagrangian (6). As we will see in this section, two conditions are required in order for the theory to propagate only two DOFs.
IV.1 The first condition
As first-class constraints, $\pi_{i}$ and $\mathcal{C}_{i}$ correspond to spatial invariance. Without breaking spatial invariance, $\pi_{i}$ and $\mathcal{C}_{i}$ will always remain first-class constraints, which have vanishing Poisson brackets with all other constraints. For simplicity, we are allowed to omit these two first-class constraints in the subsequent degeneracy analysis. We therefore define a simplified Dirac matrix by omitting the two columns and rows corresponding to $\pi_{i}$ and $\mathcal{C}_{i}$:
$\mathcal{M}^{ab}(\vec{x},\vec{y})=\begin{array}[]{|c|ccc|ccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})&\tilde{\pi}^{kl}(\vec{y})&\mathcal{C}(\vec{y})&\frac{\delta S}{\delta\phi(\vec{y})}\\ \hline\cr p(\vec{x})&0&0&0&X&X&X\\ \pi(\vec{x})&0&0&0&X&X&X\\ p^{ij}(\vec{x})&0&0&0&X&X&X\\ \hline\cr\tilde{\pi}^{ij}(\vec{x})&X&X&X&X&X&X\\ \mathcal{C}(\vec{x})&X&X&X&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&X&X&X&X&X&0\\ \hline\cr\end{array}.$ (34)
Our goal is to find conditions that eliminate one scalar-type degree of freedom. Generally, we have two options: one is to introduce extra conditions to convert second-class constraints into first-class constraints, and the other is to generate new secondary constraints. In both cases, a necessary condition is that the simplified Dirac matrix $\mathcal{M}^{ab}$ must be degenerate. The determinant of the simplified Dirac matrix $\mathcal{M}^{ab}$ is
$\det\mathcal{M}^{ab}(\vec{x},\vec{y})=\det\mathcal{B}^{ab}(\vec{x},\vec{y})\times\det\mathcal{B}^{ba}(\vec{y},\vec{x}),$ (35)
where the matrix $\mathcal{B}^{ab}$ is the lower-left corner of $\mathcal{M}^{ab}$. The degeneracy of $\mathcal{M}^{ab}$, i.e., $\det\mathcal{M}^{ab}=0$, thus implies that the sub-matrix $\mathcal{B}^{ab}$ of $\mathcal{M}^{ab}$ is degenerate, i.e., $\det\mathcal{B}^{ab}=0$.
At first glance, the determinant of the matrix $\mathcal{B}^{ab}$ may appear to be the matrix form of the first 2DOF condition. However, this is not the case. The 2DOF condition should be an equation satisfied by the action and should not contain conjugate momenta. Otherwise, we obtain a relation between the action and the conjugate momenta, which is in fact a constraint. Upon careful inspection, all elements of the matrix $\mathcal{B}^{ab}$ are independent of the conjugate momenta, except for $\mathcal{B}^{21}$, given by
$\mathcal{B}^{21}(\vec{x},\vec{y})$ $=$ $\left[\mathcal{C}(\vec{x}),p^{kl}(\vec{y})\right]$ (36)
$=$ $2\delta^{3}(\vec{x}-\vec{y})\pi^{kl}(\vec{x})-\frac{\delta^{2}S}{\delta N(\vec{x})\delta B_{kl}(\vec{y})},$
which can be rewritten as
$\mathcal{B}^{21}(\vec{x},\vec{y})\approx-\delta^{3}(\vec{x}-\vec{y})\frac{1}{N(\vec{y})}\frac{\delta S}{\delta B_{ij}(\vec{y})}+\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta N(\vec{y})}$ (37)
on the constraint surface of $\tilde{\pi}^{ij}$. In subsequent calculations, we will always use (37) as the value of $\mathcal{B}^{21}$ instead of (36). As a result, the matrix form of the first 2DOF condition becomes
$\mathcal{S}_{1}(\vec{x},\vec{y})=\det\mathcal{B}^{ab}(\vec{x},\vec{y})\approx 0,$ (38)
where
$\mathcal{B}^{ab}=\begin{array}[]{|c|ccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})\\ \hline\cr-2N(\vec{x})\tilde{\pi}^{ij}(\vec{x})&\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta\phi(\vec{y})}&N(\vec{x})\frac{\delta}{\delta N(\vec{y})}\left(\frac{1}{N(\vec{x})}\frac{\delta S}{\delta B_{ij}(\vec{x})}\right)&\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{kl}(\vec{y})}\\ -\mathcal{C}(\vec{x})&\frac{\delta^{2}S}{\delta N(\vec{x})\delta\phi(\vec{y})}&\frac{\delta^{2}S}{\delta N(\vec{x})\delta N(\vec{y})}&N(\vec{y})\frac{\delta}{\delta N(\vec{x})}\left(\frac{1}{N(\vec{y})}\frac{\delta S}{\delta B_{kl}(\vec{y})}\right)\\ \frac{\delta S}{\delta\phi(\vec{x})}&\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}&\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta N(\vec{y})}&\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta B_{kl}(\vec{y})}\\ \hline\cr\end{array}.$ (39)
In the above, we have adjusted the coefficients of $\tilde{\pi}^{ij}$ and $\mathcal{C}$ in the matrix $\mathcal{B}^{ab}$ to make it more symmetric. Since $\mathcal{B}^{ab}$ is a degenerate matrix, there must exist a null eigenvector $\mathcal{V}_{a}$ satisfying
$\int\mathrm{d}^{3}y\,\mathcal{B}^{ab}(\vec{x},\vec{y})\mathcal{V}_{b}(\vec{y})\approx 0^{a}.$ (40)
Generally, $\mathcal{V}_{a}$ takes the form
$\mathcal{V}_{a}(\vec{x})=\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)(\vec{x})=\int\mathrm{d}^{3}y\,U_{a}(\vec{x},\vec{y})V(\vec{y}).$ (41)
In (41), $V(\vec{x})$ is an arbitrary function of the spatial coordinates, and $U_{a}(\vec{x},\vec{y})$ depends on the phase space variables. One method to obtain the explicit form of the null eigenvector $\mathcal{V}_{a}$ is to calculate the adjoint matrix of $\mathcal{B}^{ab}$, as detailed in Appendix C.
Assuming that $\mathcal{B}_{2}^{*}$ does not vanish, we can use a linear combination of the null eigenvector $\mathcal{V}_{a}$ and the constraints $\tilde{\pi}^{ij}$, $\mathcal{C}$, and $\frac{\delta S}{\delta\phi}$ to define a new constraint $\mathcal{C}^{\prime}$,
$\int\mathrm{d}^{3}x\,\mathcal{C}^{\prime}(\vec{x})V(\vec{x})=\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}\tilde{\pi}^{ij}&\mathcal{C}&\frac{\delta S}{\delta\phi}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)(\vec{x})\approx 0.$ (42)
In (42), $V(\vec{x})$ is the arbitrary function from (41).
The advantage of introducing the new constraint $\mathcal{C}^{\prime}$ is that its Poisson brackets with the constraints $p$, $\pi$, and $p^{kl}$ all vanish:
$\left[\mathcal{C}^{\prime}(\vec{x}),p(\vec{y})\right]\approx\left[\mathcal{C}^{\prime}(\vec{x}),\pi(\vec{y})\right]\approx\left[\mathcal{C}^{\prime}(\vec{x}),p^{kl}(\vec{y})\right]\approx 0.$ (43)
With this new set of primary constraints, the consistency condition for the new constraint $\mathcal{C}^{\prime}$ becomes
$\frac{\mathrm{d}\mathcal{C}^{\prime}(\vec{x})}{\mathrm{d}t}$ $=\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{T}}\right]$ (44)
$=\int\mathrm{d}^{3}y\left[\mathcal{C}^{\prime}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]\lambda_{kl}(\vec{y})+\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\approx 0$
where $\lambda_{ij}$ has been fixed to zero in (26). If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\neq 0$, (44) yields a new secondary constraint $\Phi(\vec{y})$:
$\Phi(\vec{y})=\left[\mathcal{C}^{\prime}(\vec{y}),H_{C}\right]\approx 0.$ (45)
Otherwise, if no secondary constraint arises, we can still find a new first-class constraint by taking a linear combination of the existing constraints $p$, $\pi$, and $p^{ij}$. The necessary and sufficient condition for generating secondary constraints is that the consistency condition is not automatically satisfied (instead of requiring the simplified Dirac matrix to be degenerate). Therefore, after imposing the consistency condition, two scenarios may arise:
  1. 1.
    The consistency condition is not automatically satisfied, and an additional secondary constraint is needed to enforce it;
  2. 2.
    The consistency condition is automatically satisfied, and a first-class constraint can be constructed from a linear combination of existing constraints.
When we impose a specific 2DOF condition, we must check whether it violates the consistency condition and leads to additional secondary constraints. Such a newly generated constraint must be taken into account in the subsequent analysis.
IV.2 The second condition
If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]\neq 0$ in the previous section, we introduce a new constraint $\pi^{\prime}$ in the same way as $\mathcal{C}^{\prime}$,
$\int\mathrm{d}^{3}x\,\mathcal{\pi}^{\prime}(\vec{x})V(\vec{x})\equiv\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}p^{ij}&\pi&p\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}^{\prime}_{1}\right)_{ij}&\mathcal{\mathcal{V}}^{\prime}_{2}&\mathcal{\mathcal{V}}^{\prime}_{3}\end{array}\right)(\vec{x})\approx 0,$ (46)
where again $V(\vec{x})$ is the arbitrary function in (41). Similar to the constraint $\mathcal{C}^{\prime}$, the new constraint $\mathcal{\pi}^{\prime}$ has vanishing Poisson brackets:
$\left[\tilde{\pi}^{ij}(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]\approx\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]\approx\left[\frac{\delta S}{\delta\phi(\vec{x})},\mathcal{\pi}^{\prime}(\vec{y})\right]\approx 0.$ (47)
By making use of the constraints $\mathcal{\pi}^{\prime}$, $\mathcal{C}^{\prime}$ and $\Phi$, the simplified Dirac matrix becomes
=$\mathcal{M}^{\prime}{}^{ab}=\begin{array}[]{|c|ccccc|}\hline\cr[\cdot,\cdot]&\mathcal{\pi}^{\prime}(\vec{y})&\left\{p(\vec{y}),p^{kl}(\vec{y})\right\}&\mathcal{C}^{\prime}(\vec{y})&\left\{\tilde{\pi}^{kl}(\vec{y}),\frac{\delta S}{\delta\phi(\vec{y})}\right\}&\Phi(\vec{x})\\ \hline\cr\mathcal{\pi}^{\prime}(\vec{x})&0&0&0&0&X\\ \left\{p(\vec{x}),p^{kl}(\vec{x})\right\}&0&0&0&-\mathcal{C}^{ba}(\vec{y},\vec{x})&X\\ \mathcal{C}^{\prime}(\vec{x})&0&0&X&X&X\\ \left\{\tilde{\pi}^{kl}(\vec{x}),\frac{\delta S}{\delta\phi(\vec{x})}\right\}&0&\mathcal{C}^{ab}(\vec{x},\vec{y})&X&X&X\\ \Phi(\vec{x})&X&X&X&X&X\\ \hline\cr\end{array}$ (48)
Since the first 2DOF condition $\mathcal{S}_{1}(\vec{x},\vec{y})$ has been imposed, which yields a secondary constraint that reduces the dimension of the phase space by one and thus removes half a degree of freedom, we need to impose an additional 2DOF condition to reduce the number of degrees of freedom to two. This means that the simplified Dirac matrix must be degenerate again. The determinant of this matrix is proportional to $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]$, $\left[\Phi(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]$, and $\det\mathcal{C}^{ab}(\vec{x},\vec{y})$. If the determinant of this matrix is required to vanish, two types of 2DOF conditions can be obtained:
$\mathcal{S}_{2}^{(1)}(\vec{x},\vec{y})$ $=$ $\det\mathcal{C}^{ab}(\vec{x},\vec{y})\times\left[\Phi(\vec{x}),\mathcal{\pi}^{\prime}(\vec{y})\right]=\det\mathcal{D}^{ab}(\vec{x},\vec{y})\approx 0,$ (49)
$\mathcal{S}_{2}^{(2)}(\vec{x},\vec{y})$ $=$ $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]\approx 0,$ (50)
where
$\mathcal{D}^{ab}(\vec{x},\vec{y})$ $=\begin{array}[]{|c|ccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})\\ \hline\cr\tilde{\pi}^{ij}(\vec{x})&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&X&X&X\\ \Phi(\vec{x})&X&X&X\\ \hline\cr\end{array}.$ (55)
The 2DOF condition $\mathcal{S}_{2}^{(1)}$ will generate a secondary constraint if the consistency condition is not automatically satisfied. On the other hand, the 2DOF condition $\mathcal{S}_{2}^{(2)}$ clearly generates a first-class constraint.
If $\left[\mathcal{C}^{\prime}(\vec{x}),H_{\mathrm{C}}\right]=0$ in the previous section, the second 2DOF conditions become
$\mathcal{S}^{\prime}{}_{2}^{(1)}(\vec{x},\vec{y})$ $=$ $\det\mathcal{C}^{ab}(\vec{x},\vec{y})\approx 0,$ (56)
$\mathcal{S}^{\prime}{}_{2}^{(2)}(\vec{x},\vec{y})$ $=$ $\left[\mathcal{C}^{\prime}(\vec{x}),\mathcal{C}^{\prime}(\vec{y})\right]\approx 0,$ (57)
which are covered by the TTDOF conditions $\mathcal{S}_{2}^{(1)}$ and $\mathcal{S}_{2}^{(2)}$.
The space of theories satisfying the 2DOF conditions $\mathcal{S}_{1}$, $\mathcal{S}{}_{2}^{(1)}$, and $\mathcal{S}_{2}^{(2)}$ can be divided into three branches:
  1. 1.
    Linear dependence branch: If the secondary constraints are linearly dependent, the 2DOF conditions will be automatically satisfied. However, in this case, the number of degrees of freedom will not change. Due to the linear dependence, the number of secondary constraints is reduced by one, which offsets the half-degree-of-freedom reduction caused by the 2DOF conditions. For example, if the constraints $\tilde{\pi}^{ij}$, $\mathcal{C}$, and $\frac{\delta S}{\delta\phi}$ are linearly dependent, the 2DOF condition $\mathcal{S}_{1}$ is automatically satisfied. However, the number of degrees of freedom remains unchanged.
  2. 2.
    Non-physical branch: Some 2DOF conditions may render the theory trivial. For example, if the secondary constraints generated by the consistency condition still fail to satisfy the consistency condition, additional secondary constraints will be generated. These additional constraints may reduce the number of degrees of freedom to less than two, which is unphysical.
  3. 3.
    Physical branch: In this branch, we obtain a scalar-tensor theory with only two tensorial degrees of freedom.
In the following, when we refer to 2DOF conditions, we always mean those corresponding to the third physical branch.
V A concrete model of $d=2$
V.1 The Lagrangian
In [57], [58] (see also [74]), monomials of SCG have been constructed and classified according to the total number of derivatives $d$ in each monomial. As an example of our formalism, in this section we apply the two conditions to the concrete model with $d=2$. The action is given by
$\tilde{S}_{2}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},K_{ij};\mathrm{D}_{i}\right)=S_{2}\left(\phi,N,h_{ij},{}^{3}\!R_{ij},B_{ij};\mathrm{D}_{i}\right)+\int\mathrm{d}t\mathrm{d}^{3}x\frac{\delta S_{2}}{\delta B_{ij}}\left(K_{ij}-B_{ij}\right).$ (58)
As discussed earlier, we introduce the variable $B_{ij}$ to make the explicit inversion of the velocity $\dot{h}_{ij}$ in terms of the momentum possible. The equivalent action $S_{2}$ can be written as
$S_{2}=\int\mathrm{d}t\mathrm{d}^{3}x\,N\sqrt{h}\mathcal{L}_{2},$ (59)
where
$\mathcal{L}_{2}=c_{1}B_{ij}B^{ij}+c_{2}B^{2}+c_{3}R+c_{4}a_{i}a^{i}+d_{1}\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi+d_{2}a_{i}\mathrm{D}^{i}\phi.$ (60)
The coefficients $c_{i}$ and $d_{i}$ are general functions of $N$ and $\phi$. The acceleration $a_{i}$ is defined as
$a_{i}=\frac{1}{N}\mathrm{D}_{i}N.$ (61)
For later convenience, we evaluate some of the elements of $\mathcal{B}^{ab}$ as
$\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{mn}(\vec{y})}$ $=$ $2N\sqrt{h}\left(\frac{1}{2}c_{1}(\delta^{im}\delta^{jn}+\delta^{in}\delta^{jm})+c_{2}\delta^{ij}\delta^{mn}\right)(\vec{x})\delta^{3}(\vec{x}-\vec{y}),$ (62)
$N(\vec{y})\frac{\delta}{\delta N(\vec{x})}\left(\frac{1}{N(\vec{y})}\frac{\delta S_{B}}{\delta B_{ij}(\vec{y})}\right)$ $=$ $2N\sqrt{h}\left(\frac{\partial c_{1}}{\partial N}B^{ij}+\frac{\partial c_{2}}{\partial N}B\delta^{ij}\right)(\vec{x})\delta^{3}(\vec{x}-\vec{y}),$ (63)
$\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta\phi(\vec{y})}$ $=$ $2N\sqrt{h}\left(\frac{\delta c_{1}}{\delta\phi}B^{ij}+\frac{\delta c_{2}}{\delta\phi}B\delta^{ij}\right)(\vec{x})\delta^{3}(\vec{x}-\vec{y}).$ (64)
To prevent the dynamical terms in the theory from degenerating, we assume
$\det\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{mn}(\vec{y})}=\left(2N\sqrt{h}\delta^{3}(\vec{x}-\vec{y})\right)^{6}c_{1}^{6}(c_{1}+3c_{2})\neq 0,$ (65)
which implies
$c_{1}\neq 0,\qquad c_{1}+3c_{2}\neq 0.$ (66)
As long as (66) is satisfied, the non-physical branch of the 2DOF conditions is naturally removed.
Other elements of the matrix $\mathcal{B}^{ab}$ are evaluated as
$\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}$ $=$ $N\sqrt{h}\left[\left(\frac{\delta^{2}S}{\delta\phi\delta\phi}\right)_{0}+\right.$ (67)
$\quad\left(-2\frac{\partial d_{1}}{\partial\phi}\mathrm{D}^{i}\phi+\Big{(}-2N\frac{\partial d_{1}}{\partial N}-2d_{1}\Big{)}a^{i}\right)\mathrm{D}_{i}$
$\quad\left.-2d_{1}\mathrm{D}^{i}\mathrm{D}_{i}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y}),$
$\frac{\delta^{2}S}{\delta N(\vec{x})\delta N(\vec{y})}$ $=$ $N\sqrt{h}\left[\left(\frac{\delta^{2}S}{\delta N\delta N}\right)_{0}+\right.$ (68)
$\quad+\left(-\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathrm{D}^{i}\phi+\Big{(}\frac{2c_{4}}{N^{2}}-\frac{2}{N}\frac{\partial c_{4}}{\partial N}\Big{)}a^{i}\right)\mathrm{D}_{i}$
$\quad\left.-\frac{2c_{4}}{N^{2}}\mathrm{D}^{i}\mathrm{D}_{i}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y}),$
$\frac{\delta}{\delta\phi(\vec{y})}\left(\frac{\delta S}{\delta N(\vec{x})}\right)$ $=$ $N\sqrt{h}\left[\left(\frac{\delta^{2}S}{\delta\phi\delta N}\right)_{0}+\right.$ (69)
$\quad\left(\Big{(}\frac{2d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\Big{)}\mathrm{D}^{i}\phi-\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}a^{i}\right)\mathrm{D}_{i}$
$\quad\left.-\frac{d_{2}}{N}\mathrm{D}^{i}\mathrm{D}_{i}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y}),$
$\frac{\delta}{\delta N(\vec{y})}\left(\frac{\delta S}{\delta\phi(\vec{x})}\right)$ $=$ $N\sqrt{h}\left[\left(\frac{\delta^{2}S}{\delta N\delta\phi}\right)_{0}+\right.$ (70)
$\quad\left(\Big{(}-2\frac{\partial d_{1}}{\partial N}-\frac{2d_{1}}{N}\Big{)}\mathrm{D}^{i}\phi+\Big{(}\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\Big{)}a^{i}\right)\mathrm{D}_{i}$
$\quad\left.-\frac{d_{2}}{N}\mathrm{D}^{i}\mathrm{D}_{i}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y}),$
where $\left(\frac{\delta^{2}S}{\delta\phi\delta\phi}\right)_{0}$ represents the terms in $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})}$ that are independent of derivatives of the delta function. We use this notation because their explicit expressions are very complicated, and only the coefficients of the derivatives of the delta function are relevant to the subsequent analysis.
V.2 The first condition
We have already obtained the first 2DOF condition in (38). To get the specific form of the 2DOF conditions for this model, a straightforward approach is to calculate the determinant of the matrix $\mathcal{B}^{ab}$, which is, however, very complicated and not illuminating. Alternatively, we can choose an equivalent approach by assuming that a null eigenvector $\mathcal{V}_{a}(\vec{x})$ exists and find the expression of this null eigenvector. In this approach, we can determine which conditions are required to make the Dirac matrix $\mathcal{B}^{ab}$ degenerate.
We assume $U_{a}(\vec{x},\vec{y})$ in (41) is independent of the derivative of the delta function. As a result, different orders of derivatives of the delta function will be separated in the matrix $\mathcal{B}^{ab}$,
$\mathcal{B}^{ab}(\vec{x},\vec{y})=\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})+\mathcal{B}_{(1)}^{ab}(\vec{x},\vec{y})+\mathcal{B}_{(2)}^{ab}(\vec{x},\vec{y}),$ (71)
in which terms with different orders (of derivatives of the delta function) in the matrix $\mathcal{B}^{ab}$ separately satisfy
$\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(i)}^{ab}(\vec{x},\vec{y})\approx 0^{b},\quad i=0,1,2.$ (72)
Note that by observing the form of the secondary constraint $\mathcal{C}$ and $\frac{\delta S}{\delta\phi}$, (72) can only be weakly satisfied when $i=0$.
In the matrices $\mathcal{B}_{(1)}^{ab}(\vec{x},\vec{y})$ and $\mathcal{B}_{(2)}^{ab}(\vec{x},\vec{y})$, only the $\times 2$ 2 submatrices in the lower left corner are nonvanishing, which are
$\left(\begin{array}[]{cc}2\left(\frac{d_{1}}{N}+\frac{\partial d_{1}}{\partial N}-\frac{1}{N}\frac{\partial d_{2}}{\partial\phi}\right)\mathrm{D}^{i}\phi-\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}a^{i}&-\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathrm{D}^{i}\phi+2\left(\frac{c_{4}}{N^{2}}-\frac{1}{N}\frac{\partial c_{4}}{\partial N}\right)a^{i}\\ -2\frac{\partial d_{1}}{\partial\phi}\mathrm{D}^{i}\phi-2\left(N\frac{\partial d_{1}}{\partial N}+d_{1}\right)a^{i}&-2\left(\frac{\partial d_{1}}{\partial N}+\frac{d_{1}}{N}\right)\mathrm{D}^{i}\phi+\left(\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\right)a^{i}\end{array}\right)_{x}\mathrm{D}_{x^{i}}\delta^{3}\left(\vec{x}-\vec{y}\right)$ (73)
and
$\left(\begin{array}[]{cc}-\frac{d_{2}}{N}&-\frac{2c_{4}}{N^{2}}\\ -2d_{1}&-\frac{d_{2}}{N}\end{array}\right)_{x}\mathrm{D}^{x^{i}}\mathrm{D}_{x_{i}}\delta^{3}\left(\vec{x}-\vec{y}\right),$ (74)
respectively. In the above, a subscript “$x$” denotes functional dependence on $\vec{x}$. In the subsequent discussion, we omit $\vec{x}$ for brevity when the dependence on $\vec{x}$ is clear from the context.
According to (74), for the null eigenvalue equation of $\mathcal{B}_{(2)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ proportional to $\mathrm{D}^{x^{i}}\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$ yields
$d_{2}\mathcal{V}_{2}+2Nd_{1}\mathcal{V}_{3}$ $=$ 0 0,0 , (75)
$2c_{4}\mathcal{V}_{2}+Nd_{2}\mathcal{V}_{3}$ $=$ 0 0.0 . (76)
As for the null eigenvalue equation of $\mathcal{B}_{(1)}^{ab}$, the part of matrix $\mathcal{B}^{ab}$ only related to $\mathrm{D}_{x^{i}}\delta^{3}(\vec{x}-\vec{y})$ yields
$\left(\frac{2d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)\mathcal{V}_{2}-2\frac{\partial d_{1}}{\partial\phi}\mathcal{V}_{3}$ $=$ 0 0,0 , (77)
$\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2N\frac{\partial d_{1}}{\partial N}+2d_{1}\right)\mathcal{V}_{3}$ $=$ 0 0,0 , (78)
$\frac{2}{N^{2}}\frac{\partial c_{4}}{\partial\phi}\mathcal{V}_{2}-\left(2\frac{\partial d_{1}}{\partial N}+\frac{2d_{1}}{N}\right)\mathcal{V}_{3}$ $=$ 0 0,0 , (79)
$\left(\frac{2c_{4}}{N^{2}}-\frac{2}{N}\frac{\partial c_{4}}{\partial N}\right)\mathcal{V}_{2}-\left(\frac{2}{N}\frac{\partial c_{4}}{\partial\phi}-2\frac{\partial d_{2}}{\partial N}\right)\mathcal{V}_{3}$ $=$ 0 0.0 . (80)
There are four equations since the coefficients of $\mathrm{D}^{i}\phi$ and $a^{i}$ need to be proportional to the null eigenvector. We assume that $d_{1}$, $d_{2}$, and $c_{4}$ are nonvanishing. We can simply choose
$\mathcal{V}_{2}$ $=$ $2Nd_{1},$ (81)
$\mathcal{V}_{3}$ $=$ $-d_{2},$ (82)
and obtain two independent equations
$d_{2}^{2}-4d_{1}c_{4}$ $=$ 0 0,0 , (83)
$\left(\frac{d_{1}}{N}+2\frac{\partial d_{1}}{\partial N}-\frac{2}{N}\frac{\partial d_{2}}{\partial\phi}\right)2Nd_{1}+\frac{\partial d_{1}}{\partial\phi}d_{2}$ $=$ 0 0.0 . (84)
The other two equations are not independent and can be obtained by combining these two.
We have a set of special solutions
$d_{1}=\frac{A}{2N},\quad d_{2}=B,\quad c_{4}=\frac{8NB^{2}}{A},$ (85)
where $A$ and $B$ are constants. As we will see, this set of special solutions will simplify the calculations in the subsequent discussion.
When we discuss the degeneracy of the matrix $\mathcal{B}_{(0)}^{ab}$, we need to be careful to identify which parts are weakly equal to zero. To clarify this, we first discuss how secondary constraints participate in the weak equality.
Since in the Lagrangian 58 the velocity term $\dot{h}_{ij}$ (encoded in $K_{ij}$) has been linearized, each canonical momentum corresponds to a primary constraint. Then all the secondary constraints can be combined into forms that are independent of canonical momenta. For example, we make a linear combination of the constraints $\tilde{\pi}^{ij}$ and $\mathcal{C}$,
$\mathcal{C}_{1}=2\tilde{\pi}^{ij}B_{ij}-\mathcal{C}=\frac{\delta S}{\delta N}-\frac{B_{ij}}{N}\frac{\delta S}{\delta B_{ij}}.$ (86)
By combining constraints $\mathcal{C}_{1}$ and $\frac{\delta S}{\delta\phi}$, we can obtain a constraint $\mathcal{C}_{2}$,
$\mathcal{C}_{2}$ $=$ $d_{2}\frac{\delta S}{\delta\phi}-2d_{1}N\mathcal{C}_{1}$ (87)
$=$ $N\sqrt{h}\left[\left(B\frac{\partial c_{1}}{\partial\phi}-A\Big{(}-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\Big{)}\right)B_{ij}B^{ij}\right.$
$\left.+\left(B\frac{\partial c_{2}}{\partial\phi}-A\Big{(}-\frac{c_{2}}{N}+\frac{\partial c_{2}}{\partial N}\Big{)}\right)B^{2}+\left(B\frac{\partial c_{3}}{\partial\phi}-A\Big{(}\frac{c_{3}}{N}+\frac{\partial c_{3}}{\partial N}\Big{)}\right)R\right],$
which does not depend on the special solution (85) chosen before.
Since we have replaced $\mathcal{B}^{21}$ with a term independent of the canonical momentum, (38) can be equivalently written as
$\det\mathcal{B}^{ab}(\vec{x},\vec{y})=\alpha(\vec{x},\vec{y})\mathcal{C}_{2}(\vec{x})+\beta(\vec{x},\vec{y})\frac{\delta S}{\delta\phi(\vec{x})}.$ (88)
By solving equations
$\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})\approx 0^{b},\qquad b=3,4,...,8,$ (89)
we obtain
$\left(\mathcal{V}_{1}\right)_{ij}=-2d_{1}N\mathcal{V}_{ij}^{(N)}+d_{2}\mathcal{V}_{ij}^{(\phi)},$ (90)
where
$\mathcal{V}_{ij}^{(N)}$ $=$ $\frac{1}{c_{1}}\frac{\partial c_{1}}{\partial N}B_{ij}+\left(\frac{1}{c_{1}+3c_{2}}\frac{\partial c_{2}}{\partial N}-\frac{c_{2}}{(c_{1}+3c_{2})c_{1}}\frac{\partial c_{1}}{\partial N}\right)B\delta_{ij},$ (91)
$\mathcal{V}_{ij}^{(\phi)}$ $=$ $\frac{1}{c_{1}}\frac{\partial c_{1}}{\partial\phi}B_{ij}+\left(\frac{1}{c_{1}+3c_{2}}\frac{\partial c_{2}}{\partial\phi}-\frac{c_{2}}{(c_{1}+3c_{2})c_{1}}\frac{\partial c_{1}}{\partial\phi}\right)B\delta_{ij}.$ (92)
Since we have previously chosen a set of special solutions for $d_{1}$, $d_{2}$, and $c_{4}$, equation (72) with $i=0$ can be written as
$\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})=D_{1}^{b}(\vec{y})B_{ij}(\vec{y})B^{ij}(\vec{y})+D_{2}^{b}(\vec{y})B^{2}(\vec{y})+D_{3}^{b}(\vec{y})R(\vec{y}),\qquad b=1,2.$ (93)
Since there are no terms such as $a_{i}a^{i}$ in equation (93), the coefficient $\beta(\vec{x},\vec{y})$ is fixed to zero. The above formula can equivalently be written as
$\int\mathrm{d}^{3}x\,\mathcal{V}_{a}(\vec{x})\mathcal{B}_{(0)}^{ab}(\vec{x},\vec{y})=\alpha^{b}(\vec{y})\mathcal{C}_{2}(\vec{y}),\qquad b=1,2.$ (94)
In order for (94) to have a solution, the coefficients of $B_{ij}B^{ij}$, $B^{2}$, and $R$ must vanish. There will be 6 equations for $c_{1}$, $c_{2}$, and $c_{3}$. We assume that
$\frac{c_{1}}{c_{2}}=\mathrm{const}.$ (95)
This assumption simplifies the expression of $D_{2}^{b}$.
By fixing the coefficients of $B_{ij}B^{ij}$ and $R$ in (94), we obtain 4 differential equations
$A\left[\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{1}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big{(}\frac{\partial c_{1}}{\partial\phi}\Big{)}^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ 0 0,0 , (96)
$A\left[\frac{2}{N}\frac{\partial c_{1}}{\partial N}+\frac{\partial^{2}c_{1}}{\partial N^{2}}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial N}-\alpha^{2}\left(-\frac{c_{1}}{N}+\frac{\partial c_{1}}{\partial N}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{1}}{\partial\phi}+\frac{\partial^{2}c_{1}}{\partial N\partial\phi}-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial N}\frac{\partial c_{1}}{\partial\phi}-\alpha^{2}\frac{\partial c_{1}}{\partial\phi}\right)$ $=$ 0 0,0 ,
$A\left[\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{1}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ 0 0,0 ,
$A\left[\frac{2}{N}\frac{\partial c_{3}}{\partial N}+\frac{\partial^{2}c_{3}}{\partial N^{2}}-\alpha^{2}\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}\right)\right]-B\left(\frac{1}{N}\frac{\partial c_{3}}{\partial\phi}+\frac{\partial^{2}c_{3}}{\partial N\partial\phi}-\alpha^{2}\frac{\partial c_{3}}{\partial\phi}\right)$ $=$ 0 0.0 . (97)
As long as equations (95) and (96)-(97) are satisfied, the coefficient of $B^{2}$ will be automatically zero. Therefore, we do not have to discuss the differential equation corresponding to the coefficient of $B^{2}$ here.
We also choose a set of special solutions
$c_{1}$ $=$ $Nf(\phi),$ (98)
$c_{2}$ $=$ $CNf(\phi),$ (99)
$c_{3}$ $=$ $\frac{D}{N}g(\phi),$ (100)
where the coefficients $C$ and $D$ are both constants.
By plugging the expressions of $c_{1}$ and $c_{3}$ into (96)-(97), we can simplify the four differential equations as
$\frac{\partial^{2}c_{1}}{\partial\phi^{2}}-\frac{2}{c_{1}}\Big{(}\frac{\partial c_{1}}{\partial\phi}\Big{)}^{2}-\alpha^{1}\frac{\partial c_{1}}{\partial\phi}$ $=$ 0 0,0 , (101)
$\alpha^{2}\frac{\partial c_{1}}{\partial\phi}$ $=$ 0 0,0 , (102)
$\frac{\partial^{2}c_{3}}{\partial\phi^{2}}-\alpha^{1}\frac{\partial c_{3}}{\partial\phi}$ $=$ 0 0,0 , (103)
$\alpha^{2}\frac{\partial c_{3}}{\partial\phi}$ $=$ 0 0.0 . (104)
Finally, constraint $\mathcal{C}_{2}$ becomes
$\mathcal{C}_{2}$ $=BN\sqrt{h}\left(\frac{\partial c_{1}}{\partial\phi}B_{ij}B^{ij}+\frac{\partial c_{2}}{\partial\phi}B^{2}+\frac{\partial c_{3}}{\partial\phi}R\right).$ (105)
As we can see, if both $c_{1}$ and $c_{3}$ are independent of $\phi$, constraint $\mathcal{C}_{2}$ will not exist. The number of second-class constraints will be reduced by one, and the number of degrees of freedom will increase by 1/2. This is the linear-constraints branch of the 2DOF condition mentioned in the previous section, which is not the case we are interested in. Therefore, we assume that $c_{1}$ and $c_{3}$ cannot be independent of $\phi$ simultaneously. As a result, the coefficient $\alpha^{2}$ must be fixed to zero.
V.3 The second condition
According to the above discussion, after imposing the first 2DOF condition $\mathcal{S}_{1}$, we can obtain the constraint $\mathcal{C}^{\prime}$ from the null eigenvector $\mathcal{V}_{a}(\vec{x})$,
$\int\mathrm{d}^{3}x\,\mathcal{C}^{\prime}(\vec{x})V(\vec{x})\equiv\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}-2N\tilde{\pi}^{ij}&-\mathcal{C}&\frac{\delta S}{\delta\phi}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}_{1}\right)_{ij}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)(\vec{x}),$ (106)
which yields
$\mathcal{C}^{\prime}=BN\left[\sqrt{h}\left(\frac{\partial c_{1}}{\partial\phi}B^{ij}B_{ij}+\Big{(}\frac{2c_{2}}{c_{1}}\frac{\partial c_{1}}{\partial\phi}-\frac{\partial c_{2}}{\partial\phi}\Big{)}B^{2}-\frac{\partial c_{3}}{\partial\phi}R\right)-\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial\phi}\pi^{ij}B_{ij}\right].$ (107)
The Poisson brackets between the constraint $\mathcal{C}^{\prime}$ and the constraints $p$, $\pi$, and $p^{kl}$ are
$\left[\mathcal{C}^{\prime}(\vec{x}),p(\vec{y})\right]$ $=\left[-BN\frac{\partial}{\partial\phi}\left(\frac{2}{c_{1}}\frac{\partial c_{1}}{\partial\phi}\right)\tilde{\pi}^{ij}B_{ij}-\frac{\partial c_{3}}{\partial\phi\partial\phi}\left(\frac{\partial c_{3}}{\partial\phi}\right)^{-1}\mathcal{C}_{2}\right](\vec{x})\delta^{3}(\vec{x}-\vec{y})\approx 0,$ (108)
$\left[\mathcal{C}^{\prime}(\vec{x}),\pi(\vec{y})\right]=\frac{B}{N(\vec{x})}\left(\mathcal{C}^{\prime}(\vec{x})+\mathcal{C}_{2}(\vec{x})\right)\delta^{3}(\vec{x}-\vec{y})\approx 0,$ (109)
$\left[\mathcal{C}^{\prime}(\vec{x}),p^{kl}(\vec{y})\right]$ $=-\frac{BN(\vec{x})}{c_{1}(\vec{x})}\frac{\partial c_{1}(\vec{x})}{\partial\phi(\vec{x})}\tilde{\pi}^{kl}(\vec{x})\delta^{3}(\vec{x}-\vec{y})\approx 0,$ (110)
which all vanish on the constraint surface. The Poisson bracket between the constraint $\mathcal{C}^{\prime}$ and itself is generally nonzero, unless
$\frac{\partial c_{1}}{\partial\phi}=0,\quad\text{or}\quad\frac{\partial c_{3}}{\partial\phi}=0.$ (111)
As mentioned above, $c_{1}$ and $c_{3}$ cannot both be independent of $\phi$ simultaneously; otherwise, the number of degrees of freedom will increase by 1/21/21 / 2. If we require the second 2DOF condition $\mathcal{S}_{2}^{(2)}$ to hold, there are two cases.
In the first case, if we choose $\frac{\partial c_{1}}{\partial\phi}=0$, the constraint $\mathcal{C}_{2}$ becomes
$\mathcal{C}_{2}=BN\sqrt{h}\frac{\partial c_{3}}{\partial\phi}R\approx 0.$ (112)
This case is clearly non-physical, as it fixes $R$ to vanish.
On the other hand, if we choose $\frac{\partial c_{3}}{\partial\phi}=0$, the constraint $\mathcal{C}_{2}$ becomes
$\mathcal{C}_{2}$ $=BN\sqrt{h}\left(\frac{\partial c_{1}}{\partial\phi}B_{ij}B^{ij}+\frac{\partial c_{2}}{\partial\phi}B^{2}\right)\approx 0.$ (113)
In this case, we can solve for all the coefficients:
$c_{1}$ $=Nf(\phi),\quad c_{2}=CNf(\phi),\quad c_{3}=\frac{D}{N},$
$c_{4}$ $=\frac{8NB^{2}}{A},\quad d_{1}=\frac{A}{2N},\quad d_{2}=B.$ (114)
The resulting theory thus corresponds to a SCG with an auxiliary scalar field of $d=2$ that propagates two degrees of freedom. The constraint coefficients $\alpha^{i}$ in the previous differential equation are also uniquely determined:
$\alpha^{1}=-\left(\frac{\partial f(\phi)}{\partial\phi}\right)^{-1}\left(\frac{2}{f(\phi)}\Big{(}\frac{\partial f(\phi)}{\partial\phi}\Big{)}^{2}-\frac{\partial^{2}f(\phi)}{\partial\phi^{2}}\right),\quad\alpha^{2}=0.$ (115)
Now let us consider the sub-matrix $\mathcal{E}^{ab}$ of the simplified Dirac matrix $\mathcal{M}^{\prime}{}^{ab}$,
$\mathcal{E}^{ab}(\vec{x},\vec{y})=\begin{array}[]{|c|cccc|}\hline\cr[\cdot,\cdot]&p(\vec{y})&\pi(\vec{y})&p^{kl}(\vec{y})&\mathcal{C}^{\prime}(\vec{y})\\ \hline\cr\mathcal{C}^{\prime}(\vec{x})&0&0&0&0\\ \tilde{\pi}^{ij}(\vec{x})&X&X&X&X\\ \frac{\delta S}{\delta\phi(\vec{x})}&X&X&X&X\\ \Phi(\vec{x})&X&X&X&X\\ \hline\cr\end{array},$ (116)
whose determinant is obviously zero. Based on the discussion of null eigenvectors in Appendix C, we can find a null eigenvector $\mathcal{V}_{a}^{(\mathcal{E})}$ to construct a new first-class constraint $\tilde{\mathcal{C}}$,
$\int\mathrm{d}^{3}x\,\tilde{\mathcal{C}}(\vec{x})V(\vec{x})=\int\mathrm{d}^{3}x\left(\begin{array}[]{cccc}p&\pi&p^{kl}&\mathcal{C}^{\prime}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\left(\mathcal{\mathcal{V}}{}_{1}\right)_{ij}^{(\mathcal{E})}&\mathcal{\mathcal{V}}{}_{2}^{(\mathcal{E})}&\mathcal{\mathcal{V}}{}_{3}^{(\mathcal{E})}\end{array}\right)(\vec{x})\approx 0.$ (117)
To summarize, there are 24 constraints which can be divided into two classes:
$\begin{array}[]{rl}7\text{ first-class: }&\pi_{i},\quad\mathcal{C}_{i},\quad\tilde{\mathcal{C}},\\ 16\text{ second-class: }&p,\quad p^{ij},\quad\pi,\quad\tilde{\pi}^{ij},\quad\frac{\delta S}{\delta\phi},\quad\Phi.\end{array}$
The number of DOFs is thus
#$\#_{\text{dof}}$ $=\frac{1}{2}\left(2\times\#_{\mathrm{var}}-2\times\#_{1\mathrm{st}}-\#_{2\mathrm{nd}}\right)$ (118)
[…]
=2=2.= 2 .
Finally, let us discuss another 2DOF condition $\mathcal{S}_{2}^{(1)}$. By applying the consistency condition of the constraint $\mathcal{C}_{2}$, we can obtain the constraint $\Phi$,
$\Phi(\vec{x})$ $=$ $\left[\mathcal{C}^{\prime}(\vec{x}),H_{C}\right]$ (119)
$=$ $-B\frac{2N(\vec{x})}{c_{1}(\vec{x})}\frac{\partial c_{1}(\vec{x})}{\partial\phi(\vec{x})}B_{ij}(\vec{x})\frac{\delta S}{\delta h_{ij}(\vec{x})}$
$-\int d^{3}z\,BN(\vec{z})\sqrt{h}\frac{\delta}{\delta h_{ij}(\vec{z})}\left(\frac{\partial c_{1}}{\partial\phi}B^{ij}B_{ij}+\Big{(}\frac{2c_{2}}{c_{1}}\frac{\partial c_{1}}{\partial\phi}-\frac{\partial c_{2}}{\partial\phi}\Big{)}B^{2}-\frac{\partial c_{3}}{\partial\phi}R\right)(\vec{x})2N(\vec{z})B_{ij}(\vec{z}).$
Since the second-order derivative of the delta function arises when calculating the integral in the last line, the constraint $\Phi$ contains $\mathrm{D}^{i}\mathrm{D}_{j}B_{kl}$, which is proportional to
$\frac{\partial c_{3}}{\partial\phi}\mathrm{D}^{i}\mathrm{D}_{j}B_{kl}.$ (120)
Therefore, requiring the 2DOF condition $\mathcal{S}_{2}^{(1)}=\det\mathcal{D}^{ab}\approx 0$ also implies that $\frac{\partial c_{3}}{\partial\phi}$ may vanish as one of the conditions for the matrix $\mathcal{D}^{ab}$ to be degenerate, which is sufficient to satisfy the 2DOF condition $\mathcal{S}_{2}^{(2)}$.
VI Conclusion
In this work, we investigated a class of spatially covariant gravity theories with a non-dynamical scalar field. We briefly described our model in Sec. II. The general Lagrangian is given by (3), which has been shown to have two tensorial and one scalar DOFs [85]. The purpose of this work is to determine the conditions on the Lagrangian, which we dub the 2DOF conditions, under which only two degrees of freedom propagate.
A perturbative approach was taken in [86] to derive conditions such that the scalar mode is eliminated at linear order in perturbations. In this work, we employ a strict Hamiltonian constraint analysis to derive conditions such that only two DOFs propagate in the nonperturbative sense. In Sec. III, we describe the Hamiltonian formalism for the primary constraints and their consistency conditions. As expected, there are 3 DOFs in the theory if no further conditions are imposed.
In Sec. IV, by requiring the degeneracy of the Dirac matrix, we find that two conditions are required to make the theory propagate only 2 DOFs. The first condition (38) states that the matrix $\mathcal{B}^{ab}$ given in (39) must be degenerate, which results in a secondary constraint. To fully eliminate the scalar mode, a second condition is necessary. The second condition can be divided into two categories according to their different effects on the Dirac matrix. In the first case, the condition (49) generally turns a second-class constraint into a first-class constraint. In the second case, the condition (50) generates another secondary constraint $\Phi$.
We would like to emphasize that even if the Lagrangian satisfies both 2DOF conditions, the number of DOFs does not necessarily decrease. A special case is when the constraints are linearly dependent, which implies that the number of DOFs may remain unchanged. Another special case, which we refer to as the non-physical branch of the 2DOF conditions, renders the theory trivial or physically unacceptable (e.g., lacking tensor modes). One must therefore be careful when dealing with the 2DOF conditions in order to select the physical case, i.e., theories with precisely two degrees of freedom.
In Sec. V, based on the spatially covariant monomials classified in [57], [58], we consider a concrete Lagrangian (60) built from monomials of $d=2$ (i.e., with two derivatives) as an illustration of our formal analysis. Since the Dirac matrix of this model contains derivatives of the delta function, it is complicated to calculate the matrix form of the 2DOF conditions for this concrete model. Instead, we choose an equivalent but more efficient method to obtain the 2DOF conditions, which is to find the null eigenvector $\mathcal{V}_{a}$ satisfying (40). The resulting 2DOF conditions are given in (66) and (111). In particular, for this concrete model, we are able to eliminate the non-physical branch and the linear constraints branch of the 2DOF conditions. Finally, we find a set of coefficients (114), and the resulting Lagrangian corresponds to a spatially covariant gravity theory with an auxiliary scalar field that propagates only two DOFs.
Acknowledgements.
We would like to thank Zhi-Chao Wang for valuable discussions. X.G. is supported by the National Natural Science Foundation of China (NSFC) under Grants No. 12475068 and No. 11975020 and the Guangdong Basic and Applied Basic Research Foundation under Grant No. 2025A1515012977.
Appendix A Comparison with scalar-tensor theory in the spatial gauge
The spatially covariant gravity with an auxiliary scalar field was firstly proposed in [85], where the Hamiltonian analysis was performed to show that the theory propagates 3 degrees of freedom. This idea was originally motivated by generally covariant scalar-tensor theory when the scalar field possesses a spacelike gradient. After choosing the so-called “spatial gauge” as in [85], the resulting action takes a form similar to (1). However, we should note that “spatially covariant gravity with an auxiliary scalar field” and “scalar-tensor theory in the spatial gauge” are completely different theories.
First, we will show the difference between (1) and the scalar-tensor theory in the spatial gauge. Let us consider a general action of scalar-tensor theory
$S_{\mathrm{GST}}=\int\mathrm{d}^{4}x\sqrt{-g}\mathcal{L}\left(\phi;g_{ab},\varepsilon_{abcd},{}^{4}\!R_{abcd};\nabla_{a}\right),$ (121)
where the Lagrangian involves a scalar field $\phi$, the spacetime metric $g_{ab}$, the spacetime curvature tensor ${}^{4}\!R_{abcd}$ as well as their covariant derivatives. The 4-dimensional Levi-Civita tensor $\varepsilon_{abcd}$ encodes possible parity violation effects.
In a cosmological context, the scalar field $\phi$ is assumed to possess a timelike gradient so that the so-called unitary gauge with $\phi=t$ can be chosen. Here we consider the opposite situation by assuming that the scalar field possesses a spacelike gradient [^3] Similar ideas appear in “elastic inflation” [87] and “solid inflation” [88], where it is the Nambu-Goldstone boson breaking spatial diffeomorphism that plays the role of the scalar field.. Contrary to the unitary gauge which is defined by requiring $\mathrm{D}_{a}\phi=0$, we now choose a gauge in which
£$\pounds_{\bm{n}}\phi=0,$ (122)
where $n^{a}$ is the normal vector to the hypersurfaces. This can be understood as choosing hypersurfaces such that the normal vector $n^{a}$ exactly lies on the constant $\phi$ hypersurfaces. As a result, the value of $\phi$ does not change when transported along the normal vector. This choice of spatial hypersurfaces is referred to as the “spatial gauge” in [85].
The normal vector $n^{a}$ is normalized by $n^{a}n_{a}=-1$. The induced metric on the spatial hypersurfaces is, as usual,
$h_{ab}=g_{ab}+n_{a}n_{b}.$ (123)
We can then split all the 4-dimensional objects into their temporal and spatial parts. For example, the decomposition of $\nabla_{a}\phi$ is
[…]
where $\mathrm{D}_{a}$ is the covariant derivative compatible with the induced metric $h_{ab}$. In the last equality of (124) we have used the fact that £$\pounds_{\bm{n}}\phi=0$ in the spatial gauge. This is exactly the opposite of the unitary gauge, where $\mathrm{D}_{a}\phi=0$ and thus $\nabla_{a}\phi\rightarrow-n_{a}\pounds_{\bm{n}}\phi$.
After performing the 3+1 decomposition and choosing the spatial gauge, the action (121) can be written in the form
$S_{\mathrm{GST}}^{\mathrm{(s.g.)}}=\int\mathrm{d}t\mathrm{d}^{3}x\,N\sqrt{h}\,\mathcal{L}\left(\phi,N,h_{ij},{}^{3}\!R_{ij};\mathrm{D}_{i},\pounds_{\bm{n}}\right),$ (125)
where “s.g.” stands for the spatial gauge and we have fixed the spatial coordinates adapted to the spacelike hypersurfaces. At first glance, the action (125) takes the same form as (1); however, the crucial difference is that in the spatial gauge, the scalar field $\phi$ can be viewed as a time-independent but space-dependent field, which breaks spatial diffeomorphism. In other words, $\phi=\phi(\bm{x})$ in (125) cannot be viewed as a dynamical nor auxiliary variable, but is actually a “function” of spatial coordinates with fixed values. This is completely different from (1), which has spatial covariance and where $\phi$ plays the role of an auxiliary field.
Next, we show that the generally covariant version (i.e., the correspondence) of (1) is nothing but a “two-field” scalar-tensor theory. Let us take $\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi$ as an example. The covariant version is
$\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi=h^{ij}\mathrm{D}_{i}\phi\mathrm{D}_{j}\phi\rightarrow\left(g^{ab}+u^{a}u^{b}\right)\nabla_{a}\phi\nabla_{b}\phi,$ (126)
where $u_{a}=-N\nabla_{a}\Phi$ with $\Phi$ being the scalar field defining the spacelike hypersurfaces. In other words, the spacelike hypersurfaces are defined by $\Phi=\mathrm{const.}$. In (126), $N=1/\sqrt{-\left(\nabla\Phi\right)^{2}}$, which reduces to the lapse function when fixing the so-called unitary gauge with $\Phi=t$. By expanding (126) explicitly, we get
$\mathrm{D}_{i}\phi\mathrm{D}^{i}\phi\rightarrow\left(g^{ab}-\frac{\nabla^{a}\Phi\nabla^{b}\Phi}{\left(\nabla\Phi\right)^{2}}\right)\nabla_{a}\phi\nabla_{b}\phi,$ (127)
which is clearly a two-field scalar-tensor theory term.
Appendix B Constraint algebra
In this appendix, we present the explicit expressions for the Poisson brackets.
The Poisson brackets among constraints are
$\left[p(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ $=$ $\frac{1}{2N(\vec{y})}\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta B_{kl}(\vec{y})},$ (128)
$\left[p^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ $=$ $\frac{1}{2}\frac{1}{N(\vec{y})}\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta B_{kl}(\vec{y})},$ (129)
$\left[\pi(\vec{x}),\tilde{\pi}^{ij}(\vec{y})\right]$ $=$ $-\frac{1}{2}\delta^{3}(\vec{x}-\vec{y})\frac{1}{N^{2}(\vec{y})}\frac{\delta S}{\delta B_{ij}(\vec{y})}+\frac{1}{2}\frac{1}{N(\vec{y})}\frac{\delta^{2}S}{\delta N(\vec{x})\delta B_{ij}(\vec{y})},$ (130)
$\left[\tilde{\pi}^{ij}(\vec{x}),\tilde{\pi}^{kl}(\vec{y})\right]$ $=$ $\frac{1}{2N(\vec{y})}\frac{\delta^{2}S}{\delta h_{ij}(\vec{x})\delta B_{kl}(\vec{y})}-\frac{1}{2N(\vec{x})}\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta h_{kl}(\vec{y})},$ (131)
$[\pi(\vec{x}),\mathcal{C}(\vec{y})]$ $=$ $\frac{\delta^{2}S}{\delta N(\vec{x})\delta N(\vec{y})},$ (132)
${}[p(\vec{x}),\mathcal{C}(\vec{y})]$ $=$ $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta N(\vec{y})},$ (133)
$\left[p^{ij}(\vec{x}),\mathcal{C}(\vec{y})\right]$ $=$ $-2\delta^{3}(\vec{x}-\vec{y})\pi^{ij}(\vec{y})+\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta N(\vec{y})},$ (134)
$\left[\tilde{\pi}^{ij}(\vec{x}),\mathcal{C}(\vec{y})\right]$ $=$ $\frac{\delta^{2}S}{\delta h_{ij}(\vec{x})\delta N(\vec{y})}-\frac{1}{N(\vec{x})}\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta h_{kl}(\vec{y})}B_{kl}(\vec{y}),$ (135)
$\left[p(\vec{x}),\frac{\delta S}{\delta\phi(\vec{y})}\right]$ $=$ $-\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta\phi(\vec{y})},$ (136)
$\left[p^{ij}(\vec{x}),\frac{\delta S}{\delta\phi(\vec{y})}\right]$ $=$ $-\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta\phi(\vec{y})},$ (137)
$\left[\pi(\vec{x}),\frac{\delta S}{\delta\phi(\vec{y})}\right]$ $=$ $-\frac{\delta^{2}S}{\delta N(\vec{x})\delta\phi(\vec{y})},$ (138)
$\left[\tilde{\pi}^{ij}(\vec{x}),\frac{\delta S}{\delta\phi(\vec{y})}\right]$ $=$ $-\frac{\delta^{2}S}{\delta h_{ij}(\vec{x})\delta\phi(\vec{y})},$ (139)
$[\mathcal{C}(\vec{x}),\mathcal{C}(\vec{y})]$ $=$ $-\frac{\delta^{2}S}{\delta h_{ij}(\vec{y})\delta N(\vec{x})}2B_{ij}(\vec{y})+2B_{ij}(\vec{x})\frac{\delta^{2}S}{\delta h_{ij}(\vec{x})\delta N(\vec{y})},$ (140)
$\left[\frac{\delta S}{\delta\phi(\vec{x})},\mathcal{C}(\vec{y})\right]$ $=$ $\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta h_{ij}(\vec{y})}2B_{ij}(\vec{y}).$ (141)
The Poisson brackets involving the canonical Hamiltonian are
$\left[\tilde{\pi}^{ij}(\vec{x}),H_{\mathrm{C}}\right]$ $=$ $\frac{\delta S}{\delta h_{ij}(\vec{x})}-\frac{1}{N(\vec{x})}\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta B_{ij}(\vec{x})\delta h_{kl}(\vec{z})}N(\vec{z})B_{kl}(\vec{z}),$ (142)
$\left[\mathcal{C}(\vec{x}),H_{\mathrm{C}}\right]$ $=$ $2B_{ij}(\vec{x})\frac{\delta S}{\delta h_{ij}(\vec{x})}-\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta N(\vec{x})\delta h_{ij}(\vec{z})}N(\vec{z})2B_{ij}(\vec{z}),$ (143)
$\left[\frac{\delta S}{\delta\phi(\vec{x})},H_{\mathrm{C}}\right]$ $=$ $\int\mathrm{d}^{3}z\frac{\delta^{2}S}{\delta\phi(\vec{x})\delta h_{ij}(\vec{z})}N(\vec{z})2B_{ij}(\vec{z}).$ (144)
Appendix C Adjoint matrix form of null eigenvector
If we wish to find the null eigenvector of the degenerate matrix $\mathcal{B}^{ab}$, one method is to calculate the non-zero rows (or columns) of its corresponding adjoint matrix $\mathcal{B}_{ab}^{*}$, and to multiply the corresponding elements by $(-1)^{a+b}$ to build the null eigenvector. For simplicity, we assume that $\mathcal{B}^{ab}$ is a simple $\times 3$ 3 matrix and the rows and columns correspond to constraints $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$.
To verify that this form of null eigenvector is feasible, we assume that the matrix $\mathcal{B}^{ab}$ is degenerate and without loss of generality we take the first row of its adjoint matrix $\mathcal{B}_{ab}^{*}$ to form a null eigenvector
$\mathcal{V}_{a}=\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right).$ (145)
We then multiply the null eigenvector $\mathcal{V}_{a}$ by the matrix $\mathcal{B}_{ab}^{*}$,
$\mathcal{V}_{a}\mathcal{B}^{ab}=\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}^{1b}&\mathcal{B}^{2b}&\mathcal{B}^{3b}\end{array}\right)^{T},$ (146)
which yields
$\mathcal{V}_{a}\mathcal{B}^{a1}$ $=$ $\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}^{11}&\mathcal{B}^{21}&\mathcal{B}^{31}\end{array}\right)^{T}$ (149)
$=$ $\det\mathcal{B}^{ab}=0,$ (150)
$\mathcal{V}_{a}\mathcal{B}^{a2}$ $=$ $\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}^{12}&\mathcal{B}^{22}&\mathcal{B}^{32}\end{array}\right)^{T}$ (153)
$=$ $\det\left(\begin{array}[]{ccc}\mathcal{B}^{12}&\mathcal{B}^{12}&\mathcal{B}^{13}\\ \mathcal{B}^{22}&\mathcal{B}^{22}&\mathcal{B}^{23}\\ \mathcal{B}^{32}&\mathcal{B}^{32}&\mathcal{B}^{33}\end{array}\right)=0,$ (157)
$\mathcal{V}_{a}\mathcal{B}^{a3}$ $=\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}^{13}&\mathcal{B}^{23}&\mathcal{B}^{33}\end{array}\right)^{T}$ (160)
$=\det\left(\begin{array}[]{ccc}\mathcal{B}^{13}&\mathcal{B}^{12}&\mathcal{B}^{13}\\ \mathcal{B}^{23}&\mathcal{B}^{22}&\mathcal{B}^{23}\\ \mathcal{B}^{33}&\mathcal{B}^{32}&\mathcal{B}^{33}\end{array}\right)=0.$ (164)
We can combine constraints $\phi_{i}$ with the null eigenvectors to make a new constraint
$\phi^{\prime}(\vec{y})=\int\mathrm{d}^{3}x\left(\begin{array}[]{ccc}\phi_{1}&\phi_{2}&\phi_{3}\end{array}\right)(\vec{x})\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)^{T}(\vec{x},\vec{y}).$ (165)
The Poisson brackets of $\phi^{\prime}$ with constraints $\phi_{i}$ are all vanishing,
$\left[\phi_{i}(\vec{x}),\phi^{\prime}(\vec{y})\right]$$\approx$ $\int\mathrm{d}^{3}z\left[\phi_{i}(\vec{x}),\left(\begin{array}[]{ccc}\phi_{1}&\phi_{2}&\phi_{3}\end{array}\right)(\vec{z})\right]\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)^{T}(\vec{z},\vec{y})$ (168)
$=$ $\int\mathrm{d}^{3}z\left(\begin{array}[]{ccc}\mathcal{B}^{1i}&\mathcal{B}^{2i}&\mathcal{B}^{3i}\end{array}\right)\left(\begin{array}[]{ccc}\mathcal{B}_{11}^{*}&-\mathcal{B}_{21}^{*}&\mathcal{B}_{31}^{*}\end{array}\right)^{T}(\vec{z},\vec{y})$ (171)
$=$ 0 0.0 . (172)
In practical calculations, in order to avoid integrating the delta function of higher order, we need to multiply the null eigenvector by a greatest common divisor, i.e.,
$\mathcal{V}_{a}=\left(\begin{array}[]{ccc}\mathcal{\mathcal{V}}_{1}&\mathcal{\mathcal{V}}_{2}&\mathcal{\mathcal{V}}_{3}\end{array}\right)=\frac{1}{\mathrm{gcd}(\mathcal{B}_{1i}^{*},\mathcal{B}_{2i}^{*},\mathcal{B}_{3i}^{*})}\left(\begin{array}[]{ccc}\mathcal{B}_{1i}^{*}&-\mathcal{B}_{2i}^{*}&\mathcal{B}_{3i}^{*}\end{array}\right).$ (173)